### v1

@@ Abstract
Quantum Physics is a cornerstone of modern science and technology, yet a comprehensive approach to integrating it into school curricula and communicating its foundations to policymakers, industrial stakeholders, and the general public has yet to be established. In this paper, we discuss the rationale for introducing entanglement and Bell’s inequalities to a non-expert audience, and how these topics have been presented in the exhibit “Dire l’indicibile” (”Speaking the Unspeakable”), as a part of the Italian Quantum Weeks project. This initiative aims to make quantum mechanics accessible to all, bridging the gap between complex scientific principles and public understanding. Our approach meets the challenge of simplifying quantum concepts without sacrificing their core meaning, specifically avoiding the risks of oversimplification and inaccuracy. Through interactive activities – including a card game demonstration and the staging of CHSH experiments – participants explore the fundamental differences between classical and quantum probabilistic predictions. They gain insights into the significance of Bell inequality verification experiments and the implications of the 2022 Nobel Prize in Physics. Preliminary results from both informal and formal assessment sessions are encouraging, suggesting the effectiveness of this approach.
@@ 1 Introduction
The story of the Einstein, Podolsky and Rosen (EPR) paradox [1] and Bell’s inequalities [2, 3] is a paradigmatic example in the history of 20th-century physics that deserves to be communicated to students and the general public. It begins with the achievements of quantum mechanics, the best physical theory we have ever had to describe the physical world, and the problems this description entails from an epistemological point of view. One of the key points is the existence of entangled states, as described by the the axiomatic theory of quantum mechanics completed in the 1930s. The gedankenexperiment described by Einstein, Podolsky and Rosen in the paper published in 1935 [1] is precisely one such reasoning about the consequences of considering the rules of quantum mechanics as applied to the most nonclassical states we can imagine, i.e., entangled states, along with some assumptions about what we think must be true in the physical world, i.e., a definition of reality and locality. Starting from these three assumptions EPR built a paradox, whereby one of the three should be false. It is well known that the idea of EPR was that realism and locality could not be eliminated and hence that the only possibility was to consider quantum mechanics as not complete, that is not the final theory of reality, with the need to complete it with a broader new theory in which all the quantum strangeness and paradoxes would disappear. The EPR problem, posed in 1935, remained a gedankenexperiment until mid 1960s. In the meantime, extensive debate emerged regarding the consequences of eliminating one of the assumptions—realism, locality, or the completeness of quantum mechanics. The de Broglie-Bohm hidden-variables theory became the leading proposal for a causal completion of quantum mechanics [4, 5], while scientists within the Copenhagen school chose to accept quantum theory as it is, implying that realism and/or locality became not essential elements in a factual description of nature.
In 1964 John Bell reformulated the EPR problem in the framework of hidden variables and found constraints on the correlations among the outcomes of the measurements made on the two parts of a bipartite state. These constraints, called Bell’s inequalities (BI), have a very general value, just assuming realism and locality, and result in being violated if the considered bipartite state is an entangled state. Thus, the violation of Bell’s inequalities by entangled states became the crucial experiment to decide if quantum mechanics should be considered complete or not. The relevant point of BI is that they reformulated the philosophical problem set by EPR into a mathematical formula that could be tested experimentally. In the decades following Bell’s result, the problem became how to perform reliable experiments to test BI violation. It took till 2016 [6] to set up experiments that closed all the possible experimental and conceptual loopholes and could also be universally accepted, making Bell’s inequality violation probably the most scrutinized class of experiments in the history of physics.
@@
The story of the EPR paradox and BI is even more important nowadays when quantum mechanics, once a domain strictly confined to the realms of advanced physics and complex mathematics, is revolutionizing both technology and culture. The Second Quantum Revolution, driven by profound advancements in our understanding and manipulation of quantum systems, promises to transform industrial and societal perspectives at the same time. Central to this transformation is the concept of entanglement and the groundbreaking work recognized by the 2022 Nobel Prize in Physics awarded to Alain Aspect, John F. Clauser, and Anton Zeilinger for their pioneering experiments with entangled photons, establishing the violation of Bell’s inequalities and laying the foundation for quantum information science. Their work not only validated the fundamental principles of quantum mechanics but also provided a crucial boost to quantum technologies, such as quantum computing and quantum cryptography, which rely on the unique properties of entangled particles [7].
@@
We decided to make this story the core of our effort to communicate the concepts of Quantum Physics (QP) towards high-school students and general public. Though being aware of the inherent difficulty to grasp quantum concepts, the interactive exhibition, “Dire l’indicibile”, designed within the Italian Quantum Weeks project [8], strives to make QP comprehensible to everybody, using the natural language and examples from everyday experience. This initiative aims to bridge the gap between complex scientific truths and public understanding, accepting the challenge of simplifying quantum concepts without diluting their essence, avoiding the pitfalls of oversimplification and inaccuracy. The visitors of the exhibition are introduced to the basic concepts of QP, i.e., definition of quantum states, superposition, quantum measurements, non commutativity of observables, entanglement, through visual analogies and demonstrators. Explicit reference to key results, such as Feynman’s double slit and Stern-Gerlach experiments, are used to connect abstract concepts to physical evidence. Details of this path can be found in [8] and will be published elsewhere [9].
@@
In the context of education and outreach activities triggered by the ongoing Second Quantum Revolution, several efforts have been made to update the way students and general public are introduced to quantum concepts [10, 11, 12, 13]. Nevertheless, entanglement and its foundational relevance remain the most challenging concepts to communicate [14, 15]. Attempts have been made exploiting simulations and games [16, 17, 18, 19, 20], often avoiding any explicit reference to the mathematical formalism with the result of just giving suggestions and allusive results. In fact, aiming to communicate the concept of entanglement, one major challenge is the general widespread miscommunication about it, together with many other quantum concepts. For instance, a critical point in discussing entanglement is to clarify the differences between classical and quantum correlations, a distinction that is frequently overlooked or misrepresented.
@@
In this paper, we focus specifically on entanglement and Bell’s inequalities, drawing from the recent Nobel Prize-winning research. The paper is organized as follows: in Section 2 we introduce the notion of entanglement, while Section 3 provides a straightforward derivation of Bell’s inequalities in the CHSH form. Section 4 introduces our narrative approach to explaining entanglement, the EPR paradox, and Bell’s inequalities, while Sections 5 and 6 describe interactive activities demonstrating the CHSH inequality and recreating Nobel Prize-winning experiments. Section 7 presents the educational implementation of this framework and its validation. Section 8 closes the paper with some concluding remarks.
@@ 2 Entanglement
Talking about entanglement is essential as it is a fundamental phenomenon that underpins much of quantum mechanics’ non-intuitive and revolutionary aspects. Entanglement illustrates the deep and inherent connections between quantum systems, challenging classical notions of locality and realism. It plays a crucial role in both theoretical foundations and practical applications of quantum physics, making it a topic of significant interest and importance. Erwin Schroedinger, regarded entanglement as the characteristic trait of quantum mechanics, the one that enforces its entire departure from classical lines of thought [21]. This statement underscores the unique nature of entanglement, where physical systems become so deeply connected that the state of one cannot be described independently of the state of the other.
A bipartite entangled state cannot be created through local operations on the two parts alone, even if the two experimenters (Alice and Bob) are allowed to communicate. For entanglement to occur, the two systems must physically interact, resulting in a combined system that can only be described by a single wave function encompassing all possible states of the entangled pair. This means that the concept of an independent ’state’ for each subsystem loses significance, as there is no way to describe the properties of one subsystem without reference to the entire system. Consequently, observing the outcome of a measurement on one subsystem (Alice’s) allows for a precise prediction of the corresponding measurement outcome on the other (Bob’s).111 Notice that neither in the exhibition, nor in this paper, we discussed the difference between entanglement (the property of quantum state of being not preparable by local operations and classical communication) and the stronger notion of Bell-nonlocality (the property of a state to lead to violation of BI with a suitable set of measurements performed at two distant sites). Entangled states are mathematically described by a non-factorizable superpositions of different classical options for the state of multipartite systems. For a bipartite system, the overall state is entangled if it cannot be expressed as the tensor product of two states describing the two sub-systems independently.
@@
which can be viewed as a superposition of the two classical configurations for the bipartite states $|0\rangle_{A}|0\rangle_{B}$ and $|1\rangle_{A}|1\rangle_{B}$. The state in Eq. 1) cannot be written as the product of the states describing Alice’s and Bob’s sub-systems separately, that is $|\Psi^{-}_{A,B}\rangle\neq\left(a|0\rangle_{A}+b|1\rangle_{A}\right)\left(c|0\rangle_{B}+d|1\rangle_{B}\right)$ for any choice of the coefficients $a,b,c$ and $d$. We note that the structure of the entangled state in Eq. (1) implies that the output values of a measurement on system A are perfectly correlated with those of the same measurement on system B. The state remains entangled also when the basis in both subsystems are rotated by an angle $\theta$. In fact, by defining: $|0_{\theta}\rangle=\cos\theta|0\rangle+\sin\theta|1\rangle$ and $|1_{\theta}\rangle=-\sin\theta|0\rangle+\cos\theta|1\rangle$, we get
@@
In any rigorous dissemination program, it is essential to clarify the distinction between this mixed state and the “corresponding” superposition state, $\psi_{A}=\left(|0\rangle_{A}+|1\rangle_{A}\right)/\sqrt{2}$. The mixed state expresses our knowledge that the state is in either $|0\rangle_{A}$ or $|1\rangle_{A}$ whereas the superposition conveys the more challenging notion of the system existing simultaneously in both $|0\rangle_{A}$ and $|1\rangle_{A}$.
In trying to communicate the concept of entanglement in an easily-understandable way, it is quite common to dwell on the idea of perfect correlation, thus conveying the - profoundly misleading - message that Entangled states are nonclassical because they are perfectly correlated. Indeed, this is plain wrong, as also in classical physics we can produce mixed states that are perfectly correlated to each other in some variables: e.g. we could generate pairs of particles having the same color, the same size and the same weight, so that when Alice observes one property on her part of the system, she can perfectly forecast the output of Bob’s measurement of the same observable. There is no “quantum” in this result and the emphasis on the “perfect correlations” within entangled states can be misleading. What is dramatically different for entangled state is that we can also change the measurement and observe a different property which was not prepared before (e.g. the softness or the taste of the ball) and still find perfect correlations. Moreover, new observables could also be non-commuting with the others, as pointed out by EPR argument.
@@ 3 Correlation and Bell’s Inequalities
The standard way BI are usually introduced in textbooks, which is the one probed in Nobel experiments, is the CHSH formulation [22]. To understand how the CHSH inequality works, we can think of an experiment in which two people, Alice and Bob, who are in two laboratories far apart, each measure one of two parts of a system produced by some physical mechanism. The two parts of the system can be related or unrelated. Each experimenter can make measurements on their systems: Alice can choose between two possible experiments (“$A_{1}$” and “$A_{2}$”), as can Bob (“$B_{1}$” and “$B_{2}$”). The measured quantities are assumed to be dichotomic, so that the result of each individual measurement can be $+1$ or $-1$. Alice and Bob randomly and independently choose which measurement to take and record the value obtained for each repetition of the experiment. There are four possibilities for the product of the measurement values: $A_{1}\cdot B_{1}$, $A_{1}\cdot B_{2}$, $A_{2}\cdot B_{1}$, $A_{2}\cdot B_{2}$, each of which has value $+1$ or $-1$. Alice and Bob build correlation functions by calculating the average of the products of the results of all pairs of measurements taken on a large of repetitions of the experiments, namely $\langle A_{1}\cdot B_{1}\rangle$, $\langle A_{1}\cdot B_{2}\rangle$, $\langle A_{2}\cdot B_{1}\rangle$, $\langle A_{2}\cdot B_{2}\rangle$, each having a value in the interval $[-1,1]$. CHSH define the quantity
@@
that can only be $S=\pm 2$ for each pair. By taking the average of Eq. (4)
@@
we get the CHSH inequality $|\langle S\rangle|\leq 2$, where the two limiting cases correspond to perfectly correlated and anti-correlated systems, respectively. This result is general, and relies on the assumptions of realism and locality; indeed these hypothesis are implicitly assumed whenever we attribute to each $A_{i}$, $B_{i}$ (i=1,2) result a well-defined fixed value (±$\pm 1$), independent of all other measurements. BI therefore holds true for all classically correlated systems.
It is here most important to highlight that this results holds true also if hidden variables are present [2], as long as realism and locality are considered. When moving from classical to quantum systems, things change drastically. In fact, when the maximally entangled Bell’s states described in eq.1 are considered, QP predicts the violation of CHSH inequality, with $|S|$ becoming as large as $2\sqrt{(}2)$. This fundamental result is therefore in contradiction with the hypothesis of realism and locality and testing it experimentally is the route Bell indicated to probe QP completeness.
@@
    Realism: “If we can predict with certainty the result of a measurement on a system without interacting with it in any way, the measurement must correspond to a real property” [1]. A layperson’s interpretation of this idea is that the properties of physical systems (i.e., observables) hold definite values regardless of whether we measure them or not. In other words, “the moon remains in the sky even if we are not observing it” [24].
@@
    QP is incomplete. It is well known that EPR’s preference was for the third solution.
@@
    The precise interpretation of the realism and locality assumptions, particularly their non-trivial nature. While Einstein-Podolsky-Rosen were partially justified in arguing that any complete physical theory should satisfy both realist requirements (as exemplified by the ‘moon’ argument) and local causality constraints (rejecting “spooky action at a distance”), their conclusions ultimately proved incompatible with experimental evidence.
@@
    The quantitative characterization of statistical correlation and its fundamental distinction from causation. Bell’s crucial insight was recognizing that correlations provide a mathematical framework to test the EPR argument through experimental predictions. Moreover, since statistical descriptions inherently represent incomplete information, it should be clear that observed correlations between variables cannot be interpreted as causal relationships. A canonical example illustrates this principle: while sunburn incidence and ice cream consumption exhibit strong seasonal correlation, consuming ice cream does not cause sunburns. Both phenomena instead share a common causal factor – increased solar radiation during summer months.

### v2

@@ Abstract
Quantum Physics is a cornerstone of modern science and technology, yet a comprehensive approach to integrating it into school curricula and communicating its foundations to policymakers, industrial stakeholders, and the general public has yet to be established. In this paper, we discuss the rationale for introducing entanglement and Bell’s inequalities to a non-expert audience, and how these topics have been presented in the exhibit “Dire l’indicibile” (”Speaking the Unspeakable”), as a part of the Italian Quantum Weeks project. Our approach meets the challenge of simplifying quantum concepts without sacrificing their core meaning, specifically avoiding the risks of oversimplification and inaccuracy. Through interactive activities, including a card game demonstration and the VDR staging of CHSH experiments, participants explore the fundamental differences between classical and quantum probabilistic predictions. They gain insights into the significance of Bell’s inequality verification experiments and the implications of the 2022 Nobel Prize in Physics. Preliminary results from both informal and formal assessment sessions are encouraging, suggesting the effectiveness of this approach.
@@ 1 Introduction
The story of the Einstein-Podolsky-Rosen (EPR) paradox [1] and Bell’s inequalities [2, 3] stands as a paradigmatic example in the history of 20th-century physics, one that deserves communication to both students and the general public. This narrative begins with the achievements of quantum mechanics (QM), the most successful physical theory ever developed to describe the natural world, and the epistemological challenges it introduces. A central aspect of these challenges is the existence of entangled states, as formalized by the axiomatic framework of quantum mechanics completed in the 1930s.
In their seminal 1935 paper, Einstein, Podolsky, and Rosen [1] describe the alleged paradox which derives from combining the ‘natural’ assumption of reality and locality with the implications of QM rules, as applied to highly non classical states - specifically, entangled states. The consequences of their gedankenexperiment suggest therefore that at least one of these assumptions must be false.
Their conclusion was that, as realism and locality should be considered as indispensable, quantum mechanics must be incomplete and necessitating a broader theory to resolve its apparent paradoxes. The EPR problem, formulated in 1935, remained a theoretical exercise until the mid-1960s. During this period, intense debate arose over which assumption—realism, locality, or the completeness of quantum mechanics—should be abandoned. The de Broglie-Bohm hidden-variables theory emerged as a prominent attempt to provide a causal completion of quantum mechanics [4, 5], while adherents of the Copenhagen interpretation chose to accept quantum theory as it stood, effectively dismissing realism and/or locality as essential components of a factual description of nature.
In 1964, John Bell reformulated the EPR problem by assuming the existence of hidden variables and derived constraints on the correlations between measurement outcomes performed on the two parts of a bipartite quantum state. These constraints, now known as Bell’s inequalities (BI), are of profound generality, relying solely on the assumptions of realism and locality. Crucially, they are violated when the bipartite state under consideration is entangled. This violation became a decisive experimental criterion for determining whether quantum physics could be considered a complete theory. The significance of Bell’s inequalities lies in their transformation of the metaphysical problem posed by EPR into a falsifiable mathematical statement, open to experimental verification. In the decades following Bell’s work, the central challenge shifted to designing reliable experiments capable of testing the violation of these inequalities. It was not until 2016 [6] that experiments were conducted which closed all major experimental and conceptual loopholes, achieving universal acceptance. As a result, the violation of Bell’s inequalities has become one of the most rigorously tested and scrutinized classes of experiments in the history of physics.
@@
The story of the alleged EPR paradox and Bell’s inequalities (BI) holds even greater significance today, as quantum physics, once a field confined to the foundations of physics, is now revolutionizing both technology and culture. The Second Quantum Revolution, fueled by profound advancements in our understanding and control of quantum systems, promises to transform industrial and societal landscapes. At the heart of this transformation lies the concept of entanglement, a cornerstone of quantum theory. The groundbreaking contributions to this field were recognized by the 2022 Nobel Prize in Physics, awarded to Alain Aspect, John F. Clauser, and Anton Zeilinger [7] for their pioneering experiments with entangled photons, which confirmed the violation of Bell’s inequalities and laid the foundation for quantum information science. Their work not only validated the fundamental principles of quantum physics but also provided a critical impetus for the development of quantum technologies, such as quantum computing and quantum cryptography, which rely on the unique properties of entangled particles.
@@
We decided to make this story the core of our effort to communicate the concepts of Quantum Physics (QP) towards high-school students and general public. Though being aware of the inherent difficulty to grasp quantum concepts, the interactive exhibition, “Dire l’indicibile” (“Speaking the unspeakable”), designed within the Italian Quantum Weeks project [8], strives to make QP comprehensible to everybody, using the natural language and examples from everyday experience. This initiative aims to bridge the gap between complex scientific truths and public understanding, accepting the challenge of simplifying quantum concepts without diluting their essence, avoiding the pitfalls of oversimplification and inaccuracy. The visitors of the exhibition are introduced to the basic concepts of QP, i.e., definition of quantum states, superposition, quantum measurements, non commutativity of observables, entanglement, through visual analogies and demonstrators. Explicit reference to key results, such as Feynman’s double slit and Stern-Gerlach experiments, are used to connect abstract concepts to physical evidence. Details of this path can be found in [8] and will be published elsewhere [9].
@@
In the context of education and outreach activities triggered by the ongoing Second Quantum Revolution, several efforts have been made to update the way students and general public are introduced to quantum concepts [10, 11, 12, 13]. Nevertheless, entanglement and its foundational relevance remain the most challenging concepts to communicate [14, 15]. Attempts have been made exploiting simulations and games [16, 17, 18, 19, 20], often avoiding any explicit reference to the mathematical formalism with the result of just giving suggestions and allusive explanations. In fact, aiming to communicate the concept of entanglement, one major challenge is the general widespread miscommunication about it, together with many other quantum concepts. For instance, a critical point is to clarify the differences between classical and quantum correlations, a distinction that is frequently overlooked or misrepresented.
@@
In this paper, we focus specifically on entanglement and Bell’s inequalities, drawing from the recent Nobel Prize-winning research. The paper is organized as follows: in Section 2 we introduce the notion of entanglement, while Section 3 presents Bell’s inequalities in the CHSH form. Section 4 outlines our narrative approach to explaining entanglement, the EPR paradox, and Bell’s inequalities, while Sections 5 and 6 describe interactive activities demonstrating the CHSH inequality and recreating Nobel Prize-winning experiments. Section 7 presents the educational implementation of this framework and its validation. Section 8 closes the paper with some concluding remarks.
@@ 2 Entanglement
Talking about entanglement is essential as it is a fundamental phenomenon that underpins much of QP non-intuitive and revolutionary aspects. Entanglement describes the deep and inherent correlations that may be established between quantum systems, challenging classical notions of locality and realism. It plays a crucial role in both theoretical foundations and practical applications of quantum physics, making it a topic of significant interest and importance. Erwin Schroedinger, regarded entanglement as the characteristic trait of quantum mechanics, the one that enforces its entire departure from classical lines of thought [21]. This statement underscores the unique nature of entanglement, where physical systems become so deeply connected that the state of one cannot be described independently of the state of the other.
A bipartite entangled state cannot be created through local operations on the two parts alone, even if the two experimenters (say, Alice and Bob) are allowed to communicate. For entanglement to occur, the two systems must physically interact, resulting in a combined system that can only be described by a single wave function describing the properties of the entangled pair. This means that the concept of an independent ’state’ for each subsystem loses significance, as there is no way to describe the properties of one subsystem without reference to the entire system. Consequently, observing the outcome of a measurement on one subsystem (Alice’s) allows for a precise prediction of the corresponding measurement outcome on the other (Bob’s)  [^1] Notice that neither in the exhibition, nor in this paper, we discussed the difference between entanglement (the property of quantum state of being not preparable by local operations and classical communication) and the stronger notion of Bell-nonlocality (the property of a state to lead to violation of BI with a suitable set of measurements performed at two distant sites).. Entangled states are mathematically described by a non-factorizable superpositions of different classical options for the state of multipartite systems. For a bipartite system, the overall state is entangled if it cannot be expressed as the tensor product of two states describing the two sub-systems independently.
@@
which can be viewed as a superposition of the two classical configurations for the bipartite states $|0\rangle_{A}|0\rangle_{B}$ and $|1\rangle_{A}|1\rangle_{B}$. The state in Eq. 1) cannot be written as the product of the states describing Alice’s and Bob’s sub-systems separately, that is $|\Psi^{-}_{A,B}\rangle\neq\left(a|0\rangle_{A}+b|1\rangle_{A}\right)\left(c|0\rangle_{B}+d|1\rangle_{B}\right)$ for any choice of the coefficients $a,b,c$ and $d$. We note that the structure of the entangled state in Eq. (1) implies that the output values of a measurement on system A are perfectly anti-correlated with those of the same measurement on system B. The state remains entangled also when the basis in both subsystems are rotated by an angle $\theta$. In fact, by defining: $|0_{\theta}\rangle=\cos\theta|0\rangle+\sin\theta|1\rangle$ and $|1_{\theta}\rangle=-\sin\theta|0\rangle+\cos\theta|1\rangle$, we get
@@
In any rigorous dissemination program, it is essential to clarify the distinction between this mixed state and the “corresponding” superposition state, $\psi_{A}=\left(|0\rangle_{A}-|1\rangle_{A}\right)/\sqrt{2}$. The mixed state expresses our knowledge that the state is in either $|0\rangle_{A}$ or $|1\rangle_{A}$ whereas the superposition conveys the more challenging notion of the system existing simultaneously in both $|0\rangle_{A}$ and $|1\rangle_{A}$.
In trying to communicate the concept of entanglement in an easily-understandable way, it is quite common to dwell on the idea of perfect correlation (or, equivalently, anti-correlation), thus conveying the profoundly misleading message that entangled states are nonclassical because they are perfectly correlated. Indeed, this is plain wrong, as also in classical physics we can produce mixed states that are perfectly correlated to each other in some variables: e.g. we could generate pairs of particles having the same color, the same size and the same weight, so that when Alice observes one property on her part of the system, she can perfectly forecast the output of Bob’s measurement of the same observable. There is no quantumness in this result and the emphasis on the perfect correlations within entangled states can be misleading. What is dramatically different for entangled state is that we can also change the measurement and observe a different property, which was not prepared before, and still observe perfect correlations. Moreover, those new observables could also be non-commuting with the others, as pointed out by EPR argument.
@@ 3 Correlation and Bell’s Inequalities
The standard way BI are usually introduced in textbooks, which is the one probed in Nobel experiments, is the CHSH formulation [22]. To understand how the CHSH inequality works, we can think of an experiment in which two people, Alice and Bob, who are in two laboratories far apart, each measure one of two parts of a system produced by some physical mechanism. Each experimenter can make measurements on their system: Alice can choose between two possible experiments (“$A_{1}$” and “$A_{2}$”), as can Bob (“$B_{1}$” and “$B_{2}$”). The measured quantities are assumed to be dichotomic, so that the result of each individual measurement can be $+1$ or $-1$. Alice and Bob randomly and independently choose which measurement to take and record the value obtained for each repetition of the experiment. There are four possibilities for the product of the measurement values: $A_{1}\cdot B_{1}$, $A_{1}\cdot B_{2}$, $A_{2}\cdot B_{1}$, $A_{2}\cdot B_{2}$, each of which has value $+1$ or $-1$. Alice and Bob build correlation functions by calculating the average of the products of the results of all pairs of measurements taken on a large numbers of repetitions of the experiments, namely $\langle A_{1}\cdot B_{1}\rangle$, $\langle A_{1}\cdot B_{2}\rangle$, $\langle A_{2}\cdot B_{1}\rangle$, $\langle A_{2}\cdot B_{2}\rangle$, each having a value in the interval $[-1,1]$. CHSH define the quantity
@@
that can only be $S=\pm 2$ for each pair. By taking the average of Eq. (4) one obtains
@@
The CHSH inequality is a theorem that, under the assumptions of local realism and the possible existence of hidden variables, establishes the bound $|\langle S\rangle|\leq 2$ (out of a possible range of $|\langle S\rangle|\leq 4$ [23]). The two limiting cases, $\langle S\rangle=\pm 2$, correspond to systems that are perfectly correlated and anti-correlated, respectively. This result is general and rests on the assumptions of realism and locality. Specifically, these hypotheses are implicitly invoked when we assign to each measurement outcome $A_{i}$ and $B_{i}$ (for $i=1,2$) a well-defined, fixed value (±$\pm 1$), independent of all other measurements. Bell’s inequalities (BI) therefore hold true for all classically correlated systems.
As mentioned above, this results holds true also if hidden variables are present [2], as long as realism and locality are considered. When moving from classical to quantum systems, things change drastically. In fact, when the maximally entangled Bell’s states described in Eq.1 are considered, QP predicts the violation of CHSH inequality, with $|S|$ becoming as large as $2\sqrt{2}$ [24]. This fundamental result is therefore in contradiction with the hypothesis of realism and locality and testing it experimentally is the route Bell indicated to probe QP completeness.
@@
    Realism: “If we can predict with certainty the result of a measurement on a system without interacting with it in any way, the measurement must correspond to a real property” [1]. A layperson’s interpretation of this idea is that the properties of physical systems (i.e., the observables) hold definite values regardless of whether we measure them or not. In other words, “the moon remains in the sky even if we are not observing it” [26].
@@
    QP is incomplete. It is well known that EPR’s preference was for this third solution.
@@
    The fact that realism and locality assumptions can not actually be taken for granted as ‘natural’. While Einstein-Podolsky-Rosen were partially justified in arguing that any complete physical theory should satisfy both realist requirements (as exemplified by the ‘moon’ argument) and local causality constraints (rejecting “spooky action at a distance”), their conclusions ultimately proved incompatible with experimental evidence.
@@
    The quantitative characterization of statistical correlation and its fundamental distinction from causation. Bell’s crucial insight was recognizing that correlations provide a mathematical framework to test the EPR argument through experimental predictions. Moreover, since statistical descriptions inherently represent incomplete information, it should be clear that observed correlations between variables cannot be interpreted as causal relationships. A canonical example illustrates this principle: while sunburn incidence and icecream consumption exhibit strong seasonal correlation, consuming icecream does not cause sunburns. Both phenomena instead share a common causal factor – increased solar radiation during hot months.